# Variational Tail Bounds for Norms of Random Vectors and Matrices

### v3

@@ Abstract
We propose a variational tail bound for norms of random vectors under moment assumptions on their one-dimensional marginals. We also propose a simplified version of the bound that parametrizes the “aggregating” distribution in the proposed variational bound by considering a certain pushforward of the Gaussian distribution. We show that the proposed method reproduces some of the well-known bounds on norms of Gaussian random vectors, as well as a recent dimension-free concentration inequality for the spectral norm of sum of independent and identically distributed positive semidefinite matrices with sub-exponential marginals. We also obtain a similar concentration inequality for the sample covariance matrix of sub-exponential random vectors. Furthermore, we use coupling to formulate an abstraction of the proposed approach that applies more broadly.
@@ 1 Introduction
By viewing $\left\lVert X\right\rVert$} as a supremum of linear functionals indexed by the unit ball of the dual norm $\left\lVert\cdot\right\rVert$$_{*}$, variety of techniques that are developed to bound supremum of stochastic processes can be used to bound $\left\lVert X\right\rVert$}. For example, we can resort to generic chaining [Tal14] or its variants to formulate an upper bound for $\left\lVert X\right\rVert$}. If $X$ has a Gaussian distribution, Talagrand’s celebrated majorizing measures theorem [Tal87] guarantees that these upper bounds are optimal up to constant factors. However, the bounds obtained through generic chaining are usually abstract in nature and not easy to evaluate since one needs to determine an optimal (or nearly-optimal) admissible sequence of the subsets of the unit ball of $\left\lVert\cdot\right\rVert$$_{*}$. In cases where $X$ is a sum of independent and identically distributed (i.i.d.) random vectors, methods developed in empirical processes theory can be applied. Namely, using the symmetrization (see, e.g., [BLM13, Lemma 11.4]) approach, the supremum of the corresponding empirical process can be bounded in terms of the Rademacher complexity of the set of linear functionals indexed by the the unit ball of $\left\lVert\cdot\right\rVert$$_{*}$ (see [MRT19, Ch. 3], [SB14, Ch. 26], and [BLM13, Ch. 3] for the definition of Rademacher complexity and the related historical remarks).
A distinct alternative to the mentioned methods is a variational approach, commonly known as the PAC-Bayesian argument [SW97, McA98], that is primarily developed and applied to analyze the generalization error in statistical learning theory (see, e.g. [LS02, Cat07, AB07, LLS10]). More recently, PAC-Bayesian arguments are used to derive non-vacuous generalization bounds for neural networks that perfectly fit the training samples [DR17, ZVA+19]. In addition to their application in statistical learning theory, PAC-Bayesian arguments are also used in robust statistics [AC11, Cat12, CG17, Giu18] as well as non-asymptotic random matrix theory [Oli16, Zhi24]. In its most common formulation (see, e.g., [Alq24, Theorem 2.1]), to bound $\sup_{\theta\in\Theta}f_{\theta}(X)$ for a random variable $X$ and given a parameter space $\Theta$, the PAC-Bayesian argument uses the duality between the cumulant generating function and the Kullback–Leibler (KL) divergence. Specifically, the variational characterization of the KL divergence [Alq24, Lemma 2.2] guarantees that
@@
Some components of the PAC-Bayesian arguments complicate and constrain the application of this method. For example, the fundamental inequality (1) implicitly requires $f_{\theta}(X)$ to have a finite exponential moment (in a neighborhood of the origin). More importantly, we need to have an approximation for the KL divergence term in (1) that is both accurate and “calculation-friendly”. These requirements add to the subtlety of choosing a suitable reference measure $\operatorname{\mathbb{P}}_{\mathrm{ref}}$ and a suitable class of aggregating posteriors $\mathcal{P}$. For example, [Zhi24, Theorem 1 ] has established a matrix concentration inequality using the PAC-Bayesian argument in which the spectral norm is expressed as the supremum of a bilinear form, and the set $\mathcal{P}$ consists of carefully chosen truncated non-isotropic Gaussian distributions for the left and right factors of the bilinear form.
Focusing on tail bounds for the norm of random vectors (or matrices), we propose an alternative variational bound that has simpler structure and applies under fewer constraints compared to the PAC-Bayesian bounds. The role of the main components of the proposed bound are also more transparent.
@@ 2 Problem Setup and the Main Lemma
which is related to the convex geometric notions of convex cap coverings and convex floating bodies [BL88, SW90]. The central part of our results is the following elementary lemma.
@@
In fact, by straightforward changes in the derivations of Example˜2 below, we can recover, up to constant factors, the results in [Zhi24, Proposition 3] on the tail bounds for ${\left\lVert X\right\rVert}_{2}$ under the condition (5). We provide the formal statement and proof in Proposition˜1 below in Appendix˜A.
@@
holds with probability at least $1-e^{-t}$ (see, e.g., [BLM13, Example 5.7]). This special case is captured in our formulation, by choosing $N=d$ and $u_{1},\dotsc,u_{d}$ to be the canonical unit basis vectors in $\mathbb{R}^{d}$. Then, due to symmetry, we have $\operatorname{\mathbb{P}}_{0}\left(U=c_{k}u_{1}\right)=\dotsb=\operatorname{\mathbb{P}}_{0}\left(U=c_{k}u_{d}\right)=1/(2d)$. Thus
@@
Any norm can be approximated by polyhedral norms. However, using such an approximation is not always desirable because it ignores any special structure that the norm $\left\lVert\cdot\right\rVert$} might have, and may produce bounds that are not explicit in terms of the parameters that describe the law of the random vector. Lemma˜1 can still provide more explicit bounds in some special cases. As a classic example, we consider the case of bounding the $\ell_{2}$ norm of a Gaussian random vector.
Example 2 (Euclidean norm of a Gaussian random vector).
Let $X\in\mathbb{R}^{d}$ be a $\mathsf{Normal}(0,\Sigma)$ random vector for which we have
$M_{X}(u,2p)$ $\leq 2^{p}p!{\left\lVert u\right\rVert}_{\Sigma}^{2p}\,.$
We are interested in bounding ${\left\lVert X\right\rVert}_{2}$, the $\ell_{2}$-norm of $X$. Therefore, in this example, both $\mathcal{B}$ and $\mathcal{B}_{*}$ specialize to the unit $\ell_{2}$ ball.
We apply Lemma˜1 with $\operatorname{\mathbb{P}}_{0}$ being $\mathsf{Normal}(0,\kappa_{0}^{2}I)$ for some $\kappa_{0}>0$. Writing $U=\kappa_{0}G$ for a standard Gaussian random vector $G\in\mathbb{R}^{d}$, it follows from Lemma˜3 below in Appendix˜B that
$\operatorname{\mathbb{E}}_{0}\left(M_{X}(u,2p)\right)$ $\leq 2^{p}p!\,\kappa_{0}^{2p}\operatorname{\mathbb{E}}\left(\left(G^{\scriptscriptstyle\textup{{T}}}\Sigma G\right)^{p}\right)$
$\leq 2^{p}p!\,\kappa_{0}^{2p}\left(\operatorname{tr}(\Sigma)+p{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}\right)^{p}\,.$ (10)
@@
where $g\in\mathbb{R}$ is a standard Gaussian random variable, and $\Phi(\cdot)$ denotes the standard Gaussian cumulative distribution function (CDF). It follows from straightforward calculus and the inequality $2\pi<3e$ that $4\Phi(-z)-e^{-2z^{2}/3}$ is decreasing in $z\geq 0$, which implies that for all $z\geq 0$ we have [^2] Here we favored simplicity of the lower bound over its accuracy. More accurate lower bounds for the quantile function, or the Mills ratio of the standard Gaussian random variables [see e.g., [Wai19, Exercise 2.2] [BLM13, Exercise 7.8], [GU14], and references therein] can be used instead.
@@
Recalling the definition of $\nu(\operatorname{\mathbb{P}}_{0})$, we obtain
$\nu(\operatorname{\mathbb{P}}_{0})$ $\leq 2e^{2/(3\kappa_{0}^{2})}\,.$
Invoking Lemma˜1 and applying (10), we conclude that with probability at least $1-e^{-t}$, we have
${\left\lVert X\right\rVert}$ $\leq\inf_{p\geq 1,\,\kappa_{0}>0}e^{t/(2p)}\kappa_{0}\sqrt{2p}\left(\operatorname{tr}(\Sigma)+p{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}\right)^{1/2}2^{1/(2p)}e^{1/(3p\kappa_{0}^{2})}$
$=\sqrt{\frac{4e}{3}}\inf_{p\geq 1}e^{(t+\log(2))/(2p)}\left(\operatorname{tr}(\Sigma)+p{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}\right)^{1/2}$
$\leq\sqrt{\frac{8e}{3}}\left(2\operatorname{tr}(\Sigma)+\frac{1}{\log(2)}{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}t\right)^{1/2}$
$<4\sqrt{\operatorname{tr}(\Sigma)}+4\sqrt{{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}t}\,$,
where third line is obtained by evaluating the argument of the infimum at $p=1+t/\log(2)$. Therefore, up to constant factors, [^3] The provided bounds are not optimized in terms of the constant factors. Lemma˜1 recovers the tail bound for the $\ell_{2}$-norm of $X$ obtained via the Gaussian concentration inequality (see e.g., [BLM13, Theorem 5.5]), i.e., with probability at least $1-e^{-t}$, we have
@@ 3 Simplified Bounds via Pushforward of Gaussians
The tail bound of Lemma˜1 holds in rather general situations, but it is not entirely calculation-friendly, mainly due to its implicit form and the minimization with respect to $\operatorname{\mathbb{P}}_{0}$. Assuming that $X$ has sub-Gamma marginals, by characterizing $\operatorname{\mathbb{P}}_{0}$ as a certain pushforward of the standard Gaussian measure, the following theorem provides a bound that can be optimized using a few scalar parameters. In particular, one can recover the bounds in Example˜1 and Example˜2 using this theorem.
@@
$\hskip 30.00005pt\cdot\kappa_{0}\left((1-\tau){\left\lVert\Sigma\right\rVert}_{\mathrm{op}}^{1/4}\sqrt{\operatorname{tr}(\Sigma^{1/2})}+\tau\sqrt{\operatorname{tr}\Sigma}+\sqrt{2p}\sqrt{{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}}\right)\Bigg]\Bigg\}$
$\hskip 10.00002pt+2\eta_{2}p\rho_{0}\,$,
@@
${\left\lVert X\right\rVert}$ $\leq\inf_{\begin{subarray}{c}p\geq 1\\ \rho_{0}\geq 0\\ \kappa_{0}>0\end{subarray}}e^{t/(2p)}\,\Omega_{\Sigma}(p,\kappa_{0},\rho_{0})\,\overline{\nu}^{1/(2p)}_{\Sigma}(\kappa_{0},\rho_{0})\,.$ (12)
@@
$\leq T_{\geqslant}^{1/(2p)}\left(\frac{\rho_{0}}{\tau\kappa_{0}}\right)\rho_{0}{\left\lVert\Sigma\right\rVert}_{\square}^{1/2}$
$+T_{<}^{1/(4p)}\left(\frac{\rho_{0}}{\tau\kappa_{0}}\right)\left((1-\tau)\kappa_{0}{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}^{1/4}\left(\operatorname{\mathbb{E}}\left({\left\lVert G\right\rVert}_{\Sigma^{1/2}}^{4p}\right)\right)^{1/(4p)}+\tau\kappa_{0}\left(\operatorname{\mathbb{E}}\left({\left\lVert G\right\rVert}_{\Sigma}^{4p}\right)\right)^{1/(4p)}\right)\,$,
@@
In the second case, the Gaussian Poincaré inequality [^4] We apply the inequality to function $g\mapsto\left(|{\langle\nabla f_{0}(\Sigma^{1/2}g),x\rangle}|-1\right)_{+}$ which is weakly differentiable for $\kappa_{0}<\infty$. [BLM13, Theorem 3.20] yields
@@
The definition of $f_{0}$ reveals that for a certain convex function $g_{0}$ we have $f_{0}+g_{0}=\kappa_{0}{\left\lVert\cdot\right\rVert}_{\Sigma^{1/2}}^{2}/2$. It then follows from Alexandrov’s differentiability theorem (see, e.g., [EG25, Theorem 6.9], [Vil09, Theorem 14.1], and [Roc99]) that $\nabla^{2}f_{0}(\Sigma^{1/2}G)\preceq\kappa_{0}\Sigma^{-1/2}$ almost surely. Particularly, $\nabla^{2}f(\Sigma^{1/2}G)=\kappa_{0}\Sigma^{-1/2}$ whenever $\kappa_{0}{\left\lVert G\right\rVert}_{*}<\rho_{0}$. Therefore, we can write

### v4

@@ Abstract
We propose a variational tail bound for norms of random vectors under moment assumptions on their one-dimensional marginals. A simplified version of the bound that parametrizes the “aggregating distribution” using a certain pushforward of the Gaussian distribution is also provided. We apply the proposed method to reproduce some of the well-known bounds on norms of Gaussian random vectors, and also obtain dimension-free tail bounds for the Euclidean norm of random vectors with arbitrary moment profiles. Furthermore, we reproduce a dimension-free concentration inequality for sum of independent and identically distributed positive semidefinite matrices with sub-exponential marginals, and obtain a concentration inequality for the sample covariance matrix of sub-exponential random vectors. We also obtain a tail bound for the operator norm of a random matrix series whose random coefficients may have arbitrary moment profiles. Furthermore, we use coupling to formulate an abstraction of the proposed approach that applies more broadly.
@@ 1 Introduction
By viewing $\left\lVert X\right\rVert$} as a supremum of linear functionals indexed by the unit ball of the dual norm $\left\lVert\cdot\right\rVert$$_{*}$, variety of techniques that are developed to bound supremum of stochastic processes can be used to bound $\left\lVert X\right\rVert$}. For example, we can resort to generic chaining [49] or its variants to formulate an upper bound for $\left\lVert X\right\rVert$}. If $X$ has a Gaussian distribution, Talagrand’s celebrated majorizing measures theorem [48] guarantees that these upper bounds are optimal up to constant factors. However, the bounds obtained through generic chaining are usually abstract in nature and not easy to evaluate since one needs to determine an optimal (or nearly-optimal) admissible sequence of the subsets of the unit ball of $\left\lVert\cdot\right\rVert$$_{*}$. In cases where $X$ is a sum of independent and identically distributed (i.i.d.) random vectors, methods developed in empirical processes theory can be applied. Namely, using the symmetrization (see, e.g., [13, Lemma 11.4]) approach, the supremum of the corresponding empirical process can be bounded in terms of the Rademacher complexity of the set of linear functionals indexed by the the unit ball of $\left\lVert\cdot\right\rVert$$_{*}$ (see [35, Ch. 3], [46, Ch. 26], and [13, Ch. 3] for the definition of Rademacher complexity and the related historical remarks).
A distinct alternative to the mentioned methods is a variational approach, commonly known as the PAC-Bayesian argument [47, 34], that is primarily developed and applied to analyze the generalization error in statistical learning theory (see, e.g. [24, 16, 7, 31]). More recently, PAC-Bayesian arguments are used to derive non-vacuous generalization bounds for neural networks that perfectly fit the training samples [19, 57]. In addition to their application in statistical learning theory, PAC-Bayesian arguments are also used in robust statistics [8, 17, 15, 22] as well as non-asymptotic random matrix theory [38, 56]. In its most common formulation (see, e.g., [5, Theorem 2.1]), to bound $\sup_{\theta\in\Theta}f_{\theta}(X)$ for a random variable $X$ and given a parameter space $\Theta$, the PAC-Bayesian argument uses the duality between the cumulant generating function and the Kullback–Leibler (KL) divergence. Specifically, the variational characterization of the KL divergence [5, Lemma 2.2] guarantees that
@@
Some components of the PAC-Bayesian arguments complicate and constrain the application of this method. For example, the fundamental inequality (1) is useful only if $f_{\theta}(X)$ has a finite exponential moment (in a neighborhood of the origin). More importantly, we need to have an approximation for the KL divergence term in (1) that is both accurate and “calculation-friendly”. These requirements add to the subtlety of choosing a suitable reference measure $\operatorname{\mathbb{P}}_{\mathrm{ref}}$ and a suitable class of aggregating posteriors $\mathcal{P}$. For example, [56, Theorem 1 ] has established a matrix concentration inequality using the PAC-Bayesian argument in which the spectral norm is expressed as the supremum of a bilinear form, and the set $\mathcal{P}$ consists of carefully chosen truncated non-isotropic Gaussian distributions for the left and right factors of the bilinear form.
Focusing on tail bounds for the norm of random vectors (or matrices), we propose an alternative variational bound that has simpler structure and applies under fewer constraints compared to the PAC-Bayesian bounds. For example, the standard PAC-Bayesian argument does not apply under the conditions of our results in Theorems˜1, 4, and 5. The role of the main components of the proposed bound are also more transparent. Generalizations of the proposed bounds are presented in Section˜5. In particular, a PAC-Bayesian style bound is obtained as a corollary.
@@ 2 Problem Setup and the Main Lemma
which is related to the convex geometric notions of convex cap coverings and convex floating bodies [11, 45, 44]. The central part of our results is the following elementary lemma.
@@
In fact, in Corollary˜1 below we recover, up to constant factors, the results in [56, Proposition 3] on the tail bounds for ${\left\lVert X\right\rVert}_{2}$ under the condition (5).
@@
holds with probability at least $1-e^{-t}$ (see, e.g., [13, Example 5.7]). This special case is captured in our formulation, by choosing $N=d$ and $u_{1},\dotsc,u_{d}$ to be the canonical unit basis vectors in $\mathbb{R}^{d}$. Then, due to symmetry, we have $\operatorname{\mathbb{P}}_{0}\left(U=c_{k}u_{1}\right)=\dotsb=\operatorname{\mathbb{P}}_{0}\left(U=c_{k}u_{d}\right)=1/(2d)$. Thus
@@
2.1 Dimension-free tail bound for the Euclidean norm of random vectors with arbitrary moment profiles
Any norm can be approximated by polyhedral norms. However, using such an approximation is not always desirable because it ignores any special structure that the norm $\left\lVert\cdot\right\rVert$} might have, and may produce bounds that are not explicit in terms of the parameters that describe the law of the random vector. Lemma˜1 can still provide more explicit bounds in some special cases. We use Lemma˜1 to prove the following tail bound for the Euclidean norm of random vectors that have arbitrary “moment profiles”. We emphasize that the standard PAC-Bayesian argument requires exponential moments to be finite, therefore it cannot, as a standalone method, address the general situation considered by the following theorem.
Theorem 1.
Let $X\in\mathbb{R}^{d}$ be a centered random vector with covariance $\Sigma$ and a non-decreasing moment profile $h\colon{}[1,\infty]\to[0,\infty]$ that for every $u\in\mathbb{R}^{d}$ and $p\geq 2$ satisfies
$\left(\operatorname{\mathbb{E}}\left(\left|{\langle u,X\rangle}\right|^{p}\right)\right)^{1/p}$ $\leq h(p){\left\lVert u\right\rVert}_{\Sigma}\,.$ (10)
Then for any $t>0$, with probability at least $1-e^{-t}$ we have
${\left\lVert X\right\rVert}_{2}$ $\leq\inf_{p\geq 2}2p^{-1/2}h(p)\left(\operatorname{tr}\left(\Sigma\right)+\frac{p}{2}{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}\right)^{1/2}e^{(t+\log(2))/p}\,.$ (11)
Proof.
We use Lemma˜1 with $\operatorname{\mathbb{P}}_{0}$ being the law of $U=\kappa_{0}G$ for a certain $\kappa_{0}>0$ and a standard Gaussian vector $G\in\mathbb{R}^{d}$. The assumption on $X$ provides the inequality
$M_{X}(u,p)$ $\leq h^{p}(p){\left\lVert u\right\rVert}^{p}_{\Sigma}\,$,
which together with Lemma˜3 yields
$\left(\operatorname{\mathbb{E}}_{0}\left(M_{X}(U,p)\right)\right)^{1/p}$ $\leq\kappa_{0}h(p)\left(\operatorname{\mathbb{E}}\left({\left\lVert G\right\rVert}_{\Sigma}^{p}\right)\right)^{1/p}$
$\leq\kappa_{0}h(p)\left(\operatorname{tr}(\Sigma)+\frac{p}{2}{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}\right)^{1/2}\,.$
@@ 2.1 Dimension-free tail bound for the Euclidean norm of random vectors with arbitrary moment profiles
where $g\in\mathbb{R}$ is a standard Gaussian random variable, and $\Phi(\cdot)$ denotes the standard Gaussian cumulative distribution function (CDF). It follows from straightforward calculus and the inequality $2\pi<3e$ that $4\Phi(-z)-e^{-2z^{2}/3}$ is decreasing in $z\geq 0$, which implies that for all $z\geq 0$ we have [^2] Here we favored simplicity of the lower bound over its accuracy. More accurate lower bounds for the quantile function, or the Mills ratio of the standard Gaussian random variables [see e.g., [55, Exercise 2.2] [13, Exercise 7.8], [21], and references therein] can be used instead.
@@
Therefore, for all $x\in\mathcal{B}$ we have
$\operatorname{\mathbb{P}}_{0}\left(\left|{\langle U,x\rangle}\right|\geq 1\right)$ $\geq e^{-2/(3\kappa_{0}^{2})}/2\,$,
and thus
It then follows from Lemma˜1 (with the constraint $p\geq 2$ rather than $p\geq 1$) that with probability at least $1-e^{-t}$ we have
${\left\lVert X\right\rVert}_{2}$ $\leq\inf_{p\geq,\,\kappa_{0}>0}\kappa_{0}h(p)e^{2/(3p\kappa_{0}^{2})}\left(\operatorname{tr}\left(\Sigma\right)+\frac{p}{2}{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}\right)^{1/2}e^{(t+\log(2))/p}\,.$
The bound (11) follows by choosing $\kappa_{0}=2/\sqrt{3p}$ and using the fact that $\sqrt{e/3}<1$. ∎
The following corollary provides more explicit bounds for random vectors that have sub-Gaussian or sub-exponential marginals.
Corollary 1.
Under the conditions of Theorem˜1, if $X$ has sub-Gaussian marginals, in the sense that $h(p)=\eta p^{1/2}$ in (10) for some $\eta\geq 1$, then for any $t>0$, with probability at least $1-e^{-t}$ we have
${\left\lVert X\right\rVert}_{2}$ $\leq 6\eta\left(\operatorname{tr}\left(\Sigma\right)+t{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}\right)^{1/2}\,.$ (13)
Similarly, if $X$ has sub-exponential marginals, i.e., $h(p)=\eta p$ for some $\eta\geq 1$, then for any $t\geq 1$, with probability at least $1-e^{-t}$ we have
${\left\lVert X\right\rVert}_{2}$ $\leq 4\sqrt{e}\eta\left(\sqrt{t\operatorname{tr}\left(\Sigma\right)}+t{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}^{1/2}\right)\,.$ (14)
Proof.
Let $r=\operatorname{tr}(\Sigma)/{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}$. Applying Theorem˜1 in the sub-Gaussian case, with probability at least $1-e^{-t}$ we have
${\left\lVert X\right\rVert}_{2}$ $\leq\inf_{p\geq 2}2\eta{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}^{1/2}\left(r+\frac{p}{2}\right)^{1/2}e^{(t+\log(2))/p}\,$,
which by choosing $p=2(t+\log(2)+\sqrt{r(t+\log(2))})\geq 2(\log(2)+\sqrt{r\log(2)})>2$ yields
${\left\lVert X\right\rVert}_{2}$ $\leq 2\eta{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}^{1/2}\left(r+t+\log(2)+\sqrt{r(t+\log(2))}\right)^{1/2}\sqrt{e}$
$\leq\sqrt{6e}\eta{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}^{1/2}\left(r+t+\log(2)\right)^{1/2}$
$\leq 6\eta{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}^{1/2}\left(r+t\right)^{1/2}$
$=6\eta\left(\operatorname{tr}\left(\Sigma\right)+t{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}\right)^{1/2}\,$,
where in the second line we used $2\sqrt{r(t+\log(2))}\leq r+t+\log(2)$, and in the third line we used the facts that $r\geq 1$ and $e(1+\log(2))<6$.
In the sub-exponential case, Theorem˜1 guarantees, with probability at least $1-e^{-t}$, that
${\left\lVert X\right\rVert}_{2}$ $\leq\inf_{p\geq 2}2\eta{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}^{1/2}\left(pr+\frac{p^{2}}{2}\right)^{1/2}e^{(t+\log(2))/p}\,.$
Since $t\geq 1$, choosing $p=2t$ guarantees that $e^{(t+\log(2))/p}\leq\sqrt{2e}$ and thus
${\left\lVert X\right\rVert}_{2}$ $\leq 4\sqrt{e}\eta{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}^{1/2}\left(tr+t^{2}\right)^{1/2}$
$\leq 4\sqrt{e}\eta\left(\sqrt{t\operatorname{tr}\left(\Sigma\right)}+t{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}^{1/2}\right)\,$,
where the second line follows from the sub-additivity of the square root. ∎
For random vectors with sub-Gaussian marginals, the tail bounds of the form (13) are well-known and can be derived using more standard methods (see, e.g., [53, Exercise 6.3.5]). If $X$ is a centered Gaussian, the Gaussian concentration inequality [13, Theorem 5.5] provides a better approximation for the tail behavior of ${\left\lVert X\right\rVert}_{2}$ and guarantees that with probability at least $1-e^{-t}$ we have
@@
Therefore, with $h(p)=\sqrt{2p/\pi}$ being a valid moment profile for the Gaussian random vector $X$, (13) recovers the inequality above up to a factor of $6\sqrt{2/\pi}<5$.
The tail bound (14), up to the constant factors, is equivalent to the tail bound due to [56, Proposition 3 ] for random vectors with sub-exponential marginals. As pointed out in [56], this bound applies to log-concave measures, including uniform measures on convex bodies, and as such it improves on a similar result due to [4]. Also, it can be viewed as a dimension-free specialization of a general result of [25] to the case of the Euclidean norm of sub-exponential random vectors.
@@ 3 Simplified Bounds via Pushforward of Gaussians
The tail bound of Lemma˜1 holds in rather general situations, but it is not entirely calculation-friendly, mainly due to its implicit form and the minimization with respect to $\operatorname{\mathbb{P}}_{0}$. Assuming that $X$ has sub-Gamma marginals, by characterizing $\operatorname{\mathbb{P}}_{0}$ as a certain pushforward of the standard Gaussian measure, the following theorem provides a bound that can be optimized using a few scalar parameters. In particular, one can recover the tail bound in Example˜1, and (13) of Corollary˜1 using this theorem. The analysis for the latter is provided in Appendix˜A.
@@
$\hskip 30.00005pt\cdot\kappa_{0}\left((1-\tau){\left\lVert\Sigma\right\rVert}_{\mathrm{op}}^{1/4}\sqrt{\operatorname{tr}(\Sigma^{1/2})}+\tau\sqrt{\operatorname{tr}\Sigma}+\sqrt{2p}\sqrt{{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}}\right)\Bigg]\Bigg\}+2\eta_{2}p\rho_{0}\,$,
@@
${\left\lVert X\right\rVert}$ $\leq\inf\left\{e^{t/(2p)}\,\Omega_{\Sigma}(p,\kappa_{0},\rho_{0})\,\overline{\nu}^{1/(2p)}_{\Sigma}(\kappa_{0},\rho_{0})\colon{}p\geq 1,\,\rho_{0}\geq 0,\,\kappa_{0}>0\right\}\,.$ (15)
@@
$\leq T_{\geqslant}^{1/(2p)}\left(\frac{\rho_{0}}{\tau\kappa_{0}}\right)\rho_{0}{\left\lVert\Sigma\right\rVert}_{\square}^{1/2}+$
$T_{<}^{1/(4p)}\left(\frac{\rho_{0}}{\tau\kappa_{0}}\right)\left((1-\tau)\kappa_{0}{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}^{1/4}\left(\operatorname{\mathbb{E}}\left({\left\lVert G\right\rVert}_{\Sigma^{1/2}}^{4p}\right)\right)^{1/(4p)}+\tau\kappa_{0}\left(\operatorname{\mathbb{E}}\left({\left\lVert G\right\rVert}_{\Sigma}^{4p}\right)\right)^{1/(4p)}\right)\,$,
@@
In the second case, the Gaussian Poincaré inequality [^3] We apply the inequality to function $g\mapsto\left(|{\langle\nabla f_{0}(\Sigma^{1/2}g),x\rangle}|-1\right)_{+}$ which is weakly differentiable for $\kappa_{0}<\infty$. [13, Theorem 3.20] yields
@@
The definition of $f_{0}$ reveals that for a certain convex function $g_{0}$ we have $f_{0}+g_{0}=\kappa_{0}{\left\lVert\cdot\right\rVert}_{\Sigma^{1/2}}^{2}/2$. It then follows from Alexandrov’s differentiability theorem (see, e.g., [20, Theorem 6.9], [54, Theorems 14.25 and 14.1], and [43]) that $\nabla^{2}f_{0}(\Sigma^{1/2}G)\preceq\kappa_{0}\Sigma^{-1/2}$ almost surely. Particularly, $\nabla^{2}f(\Sigma^{1/2}G)=\kappa_{0}\Sigma^{-1/2}$ whenever $\kappa_{0}{\left\lVert G\right\rVert}_{*}<\rho_{0}$. Therefore, we can write